OpenStudy (anonymous):

What is the solution to the equation 36^4c = sqrt of 6^(8c - 6) ?

OpenStudy (anonymous):

6^(2*4c) = 6^((1/2)*(8c-6)) as bases are both 6 8c = 4c - 3 4c = -3 c = -3/4

OpenStudy (anonymous):

thans

OpenStudy (anonymous):

ok